a metallic sphere of radius R is kept in the path of a parallel beam of light of intensity I.the force exerted by beam on the sphere (where c=velocity of light & metallic sphere absorbed all light.


 

Grade:12

1 Answers

Aman Bansal
13 years ago

dear sushant raj,

momentum p= u/c, u=energy & c= velocity of light,

 I= U/A , a= area

for the sake of simplicity we are taking area A to be  ΠR2

therefore u= I*A=I*pi *R2

=> p = I*pi *R2/c

and f=dp/dt

considering that light strikes for a small time interval of delta t

therefore F= I*pi *R2/c*deltat

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar

aman bansal